A car of mass 1200kg traveling westward at 30. M/s is slowed to a stop in a distance of 50. M by the cars brakes. What was the breaking force?
1
Expert's answer
2019-10-18T10:09:46-0400

The magnitude of the acceleration of a car


"a=\frac{v^2}{2d}=\frac{(30\:\rm m/s)^2}{2\times 50\:\rm m}=9\:\rm m/s^2"

The magnitude of the braking force


"F=ma=1200\:\rm rg\times 9\: m/s^2=10800\: N"

The direction of the braking force is toward to the East.